# Strain stiffening due to stretching of entangled particles in random packings of granular materials

### v1

@@ III.4 Stretching of links
The relative size of the two peaks in Fig. 7a indicates how many links are compressed or stretched. To distinguish the peaks from the background, the distribution $p_{d}$ is fit to the sum of three Gaussian functions – the left peak for compressed chains, the right peak for fully extended chains, and a broad middle Gaussian for the background. The fraction of fully extended links $p_{s}$ is calculated as the integral of the Gaussian fit to the rightmost peak of the probability distribution $p_{d}(d/a)$ and shown in Fig. 7b for each $N$ and $\gamma$. This indicates a significantly increased fraction of stretched links when the packing is strained, and the change in the fraction of extended links is similar for different $N$.
While the integral of the rightmost peak of Fig. 7a indicates the fraction of links that are fully extended so that they may stretch, the average strain $\gamma_{p}$ of those links that are stretched can be determined from the change in position of the rightmost peak. Comparing the positions $d_{R}$ of the rightmost peaks in Fig. 7a indicates an additional link strain $\gamma_{p}=[d_{R}(\gamma=0.2)-d_{R}(\gamma=0)]/d_{R}(\gamma=0)=0.016\pm 0.002$ for $N=20$. This average strain corresponds to an average tensile load of 44 N on each stretched link (see Fig. 17 in Appendix B). This is a significant fraction of the average total load of 1400 N on strong strain stiffening samples at $\gamma=0.2$, which suggests that stretched links could support a significant part of the load in strong strain stiffening samples  [^1] We could not calculate the strain for $N=10$ because there was not a clearly resolvable peak in $p_{d}$ at $\gamma=0$, most likely due to that packing being initially very loose (see Fig. 21 in Appendix D). In contrast, for $N=5$, we measure a $\gamma_{p}=0.001\pm 0.002$, consistent with zero and much smaller than the value for $N=20$. This confirms that there is no resolvable stretching of links for the chains that are too short to have system-filling networks of entanglements at $N=5$.
@@
The autocorrelation of center-to-center distances $d$ is shown as a function of the separation distance $\Delta m$ along a chain in units of beads in Fig. 8. There is a significant variation of these correlation lengths among the packings at $\gamma=0$, suggesting the initial packing structure might affect this. An exponential fit to both $N=10$ and $N=20$ data at $\gamma=0.2$ for $m\leq 5$ yields a characteristic correlation length of $\lambda=1.1\pm 0.1$ beads. This short correlation length suggests that entangled chains do not consist of a sequence where every link in the same chain is fully extended, i.e. with the force chain following the links of the chain. Rather, the stretched links can still act as extra constraints on the packing where large forces can be supported, but these forces may be transmitted along frictional contacts with other beads instead of just along the links of a single chain.
@@ III.5 Prediction of the stress-strain curve
Now that we have established that strain stiffening occurs along with stretching of links, but only in packings with system-filling clusters of entanglements, we desire a predictive model of how that leads to strong strain stiffening. Since the stiffness of the chains is much higher than that of the elastic boundary, without constraints that can provide support across the entire system, the stiffness of the overall system would be limited by the weaker stiffness of the boundary. Importantly, system-filling clusters of entanglements allow for a possible way to provide additional internal constraints that prevent failure by shear banding and provide an effective replacement for confinement at the boundary. With such constraints, the strength can exceed the confining pressure at the boundary by stretching of the chains. While the strength can originate from the stretching of chains, the constraints of entanglement may allow forces to be transmitted along a contact network that includes the entire system when connected by system-filling clusters of entanglements.
To calculate the stress-strain relation we use a mean-field mode where we assume the average stress $\sigma$ measured on the endplate in response to triaxial load ultimately comes from and is equal to a volume average of the magnitude of the stress on the surface of particles. This matching of magnitudes regardless of the direction of forces is reasonable if those contact forces are redirected randomly along a force network. To relate the global stress $\sigma$ to the particle stress $\sigma_{p}$ and particle strain $\gamma_{p}$, we write the derivative of the global stress-strain relation using the chain rule from calculus
@@
The angle brackets represent a mean-field average to be calculated as a volume average over a typical local structure. The effective extensional modulus of the chain $\partial\sigma_{p}/\partial\gamma_{p}\equiv E_{p}$ is calculated as the force pulling on the chain divided by the average cross-sectional area of the chain and the strain of the chain (see Appendix B). $\langle\partial\sigma/\partial\sigma_{p}\rangle$ relates the average stress over the entire volume to the average stress on a chain, so the volume-averaged stress corresponds to the fraction of the volume that is being stretched, and is calculated as the packing fraction $\phi$ times the probability of stretched links $p_{s}$. $\langle\partial\gamma_{p}/\partial\gamma\rangle$ then relates the particle strain to the applied strain $\gamma$ averaged over links that are stretched. Equation 1 can now be rewritten as
@@
Figure 9: (color online) Fitted curvature $a_{2}$ of the stress-strain data from Fig. 3, including the 30% run-to-run variation in the uncertainty. Dotted line: model prediction for $a_{2}$ due to system-filling clusters of entanglements. The model prediction is consistent with each curvature for $N\geq 9$ where strong strain stiffening is observed.
For comparison to the model, fits of a quadratic function $\sigma=a_{2}\gamma^{2}$ to each measured $\sigma(\gamma)$ in Fig. 3 up to $\gamma=0.2$ are shown in Fig. 9. For each $N\geq 9$ where strong strain stiffening is found, the predicted curvature $a_{2}$ based on structural measurements with a 25% uncertainty is consistent with each fit including the 30% run-to-run variation on the measurements. Fitting the curve at $N=10000$ up to a larger strain of 0.45 results in less than a 2% change in $a_{2}$, confirming the scaling is similar over a wider parameter range. Since the curvature $a_{2}$ of the measured $\sigma(\gamma)$ for $N\geq 9$ is consistent with the model prediction, the strength of strain stiffening can be attributed to the stretching of links, and the quadratic shape of strain stiffening can be attributed to a linear increase in the fraction of links being stretched $p_{s}$ with strain.
@@
To confirm whether the measured stress-strain curves $\sigma(\gamma)$ follow the predicted quadratic curvature, we replot each $\sigma(\gamma)$ from Fig. 3 on a log-log scale in Fig. 10. The strong strain-stiffening curves with $N\geq 9$ are seen to be parallel to the predicted quadratic curve with slope 2 for strains $\gamma\stackrel{{\scriptstyle>}}{{{}_{\sim}}}0.04$, showing that the quadratic prediction correctly describes these data where strong strain stiffening occurs. For small strain $\gamma\stackrel{{\scriptstyle<}}{{{}_{\sim}}}0.01$, these curves have a nearly linear stress-strain relation, however in this range the stress is consistently less than 10 kPa, which not not reach significantly higher strength than packings of individual beads ($N=1$, see inset of Fig. 2), so it does not characterize the much higher strength which makes the strong strain stiffening of interest. A transition from strain softening to strong strain stiffening can be seen for the curves at smaller $N$, which gradually approach a slope of 2 with increasing $N$, although a big jump in the magnitude of the stress remains between $N=8$ and 9.
In Eq. 3, we could have assumed a constant added to $p_{s}$ corresponding to the fraction of links fully stretched at zero strain, which could be as large as 0.09 based on the $N=20$ data in Fig. 7b. This would have added a linear term to the stress-strain relation, corresponding to the nearly linear regime in Fig. 10 for small strain. To test if this nearly linear stress-strain regime is related to $p_{s}$ at zero strain, we added a linear term to the polynomial fit $\sigma=a_{1}\gamma+a_{2}\gamma^{2}$ to the data. Fits of the stress-strain relation yield an average $a_{1}=0.3\pm 0.2$ MPa, which would correspond to a probability of stretched links of $p_{s}=a_{1}/E_{p}\phi\langle\partial\gamma_{p}/\partial\gamma\rangle=0.004\pm 0.003$. This is much smaller than the number of links that are fully extended at $\gamma=0$ as seen in Fig. 7, suggesting that most of these initially extended links are not stretched enough for the link stiffness to contribute to the packing strength. A negligible initial stretching of links should be expected as long as the initial packing is done without a significant loading.
@@ IV Generalizing the strain stiffening model to staples
Equation 2 represents a simple model which attributes the large stress in entangled structures to the deformation of particles, with strain stiffening due to an increasing number of deformed particles with strain. This concept could be extended to a variety of entangled particle systems, not just limited to chains. To test the generality of this model for strain stiffening we consider a very different particle shape: rigid U-shaped staples, which can also entangle. Staples have a few features that are helpful for testing the stress-strain model beyond simply having another shape as a data point. Since the particles are rigid, geometric calculations can be made to determine the free entangled area and probability of deformation $p_{s}$ without need for measurements. Because the particles deform plastically, we can measure the bend angles of staple arms after compression of packings to infer information about entanglement and extensional forces between particles without the need for x-ray tomography, and thus obtain data more easily, which is helpful for example to test the strain dependence of factors such as $p_{s}$ and $\langle\partial\gamma_{p}/\partial\gamma\rangle$.
@@ IV.1 Strain stiffening
Figure 11: (color online) Another example of strain stiffening in randomly packed particles that can entangle – in this case rigid U-shaped staples. The black curve indicates the mean and the gray band indicates the standard deviation of three measurements. The red dotted line shows the model prediction of Eq. 6 due to stretching of staples, which is consistent with the measured curvature within the 36% uncertainty on the model.
@@
The staples enclose an area that corresponds to a free enclosed area $A$ of 80 times the arm or backbone cross-section $wt$ (see Fig. 1c), which is large compared to the longest chains we measured in Fig. 6, so a large number of entanglements is expected. Likewise, simulations of staple packings for the aspect ratio $L_{x}/L_{y}=0.37$ produce an average of 6 entanglements per particle Gravish et al. (2012), more than enough for system-filling clusters according to the Erdös-Rényi model. Our observation of strain-stiffening staple packings is consistent with our observation for chains in Sec. III.2 that average effective free enclosed areas that are large enough to entangle more than one particle result to strain-stiffening behavior. Since the number of entanglements as a function of arm length has already been studied Gravish et al. (2012), here we focus on identifying the source of stress and extending the model for the stress-strain relation to staples, which is less straightforward.
@@ IV.2 Particle bending observations
To identify the source of stress, we look at plastically deformed staples after triaxial compression of a packing. Figure 12 shows a random sampling of 102 plastically bent staples that were poured out and sorted after a triaxial compression test to a strain of 47%. The bends can be categorized based on where contact forces would have to be to produce the bends. 20% of staples showed no resolvable plastic deformation, indicating that the stresses on those particles were below the yield stress for plastic deformation. 45% of staples have the arms bent outward, suggesting they were pulled apart due to extensional loadings on both arms. Another 19% of staples had arms bent outward with an additional inward bend in the backbone. This bend of the backbone could be due to bending from tensile forces on the arms if the contact points are offset from the backbone, or the backbone could be bent due to an additional contact pushing directly on the backbone along with tensional forces on each arm. Combined, this is a total of 64% of the staples that are clearly under some sort of tension. 17% of staples had deformations that indicate a mixture of extensional forces on one arm and compression loadings on the other arm, corresponding to a more complex force balance requiring multiple contact points for mechanical equilibrium which are not apparent from the plastic deformation alone. Notably, no staples were bent in such a way to indicate compressive forces on both arms, even though the sample overall was under compression. Overall, the vast majority of deformed staples experienced extensional forces on the arms, indicating extensional forces from entangling were dominant. This is consistent with previous simulations of packings that found tensile stress on S-shaped and U-shaped particles Karapiperis et al. (2022), as well as expectations based on the observations of chains in Sec. III.4 that tension on particles plays an important role in strain stiffening.
@@ IV.3 Model for stress
The model of Eq. 2 for the strength of entangled packings expresses the slope of the stress-strain curve in terms of the probability of particles being stretched $p_{s}$, and the ratio of local stain to global strain $\langle\partial\gamma_{p}/\partial\gamma\rangle$. Since the plastic deformation of staples is more easily expressed in terms of a permanent plastic deformation angle of the arm $\theta_{p}$ instead of the linear strain of a chain $\gamma_{p}$, we rewrite the chain rule of Eq. 1 in terms of the arm bend angle $\theta_{p}$. While the load on particles is applied in the direction of tension, the bending in Fig. 12 suggests that the internal stress is mainly bending stress at the corners of the staples, and perhaps in the backbone. For extensional forces $F_{p}$ applied to the arms at a lever arm distance $l$ away from the corners of the staple, the average stress at the surface of the staples due to bending is approximately $\sigma_{p}=6lF_{p}L_{y}/wt^{2}(L_{y}+2L_{x})$ (see Appendix F for derivation of relationship between normal stress and force). Thus, an appropriate application of Eq. 1 and 2 to staples is

### v2

@@ III.4 Stretching of links
The relative sizes of the two peaks in Fig. 7a indicates how many links are compressed or fully extended. To distinguish the peaks from the background, the distribution $p_{d}$ is fit to the sum of three Gaussian functions – the left peak for compressed chains, the right peak for fully extended chains, and a broad middle Gaussian for the background. The fraction of fully extended links $p_{s}$ is calculated as the integral of the Gaussian fit to the rightmost peak of the probability distribution $p_{d}(d/a)$ and shown in Fig. 7b for each $N$ and $\gamma$. This indicates a significantly increased fraction of stretched links when the packing is strained, and the change in the fraction of extended links is similar for different $N$.
While the integral of the rightmost peak of Fig. 7a indicates the fraction of links that are fully extended so that they may stretch, the average strain $\gamma_{p}$ of those links that are stretched can be determined from the change in position of the rightmost peak. Comparing the positions $d_{R}$ of the rightmost peaks in Fig. 7a indicates an additional link strain $\gamma_{p}=[d_{R}(\gamma=0.2)-d_{R}(\gamma=0)]/d_{R}(\gamma=0)=0.016\pm 0.002$ for $N=20$. This average strain corresponds to an average tensile load of $44\pm 6$ N on each stretched link (see Fig. 17 in Appendix B). This is a significant fraction of the average total load of 1400 N on strong strain stiffening samples at $\gamma=0.2$, which suggests that stretched links could support a significant part of the load in strong strain stiffening samples  [^1] We could not calculate the strain for $N=10$ because there was not a clearly resolvable peak in $p_{d}$ at $\gamma=0$, most likely due to that packing being initially very loose (see Fig. 21 in Appendix D). In contrast, for $N=5$, we measure a $\gamma_{p}=0.001\pm 0.002$, consistent with zero and much smaller than the value for $N=20$. This confirms that there is no resolvable stretching of links for the chains that are too short to have system-filling networks of entanglements at $N=5$.
@@
The autocorrelation of center-to-center distances $d$ is shown as a function of the separation distance $\Delta m$ along a chain in units of beads in Fig. 8. There is a significant variation of correlation lengths among the packings at $\gamma=0$, suggesting the initial packing structure might affect this. An exponential fit to both $N=10$ and $N=20$ data at $\gamma=0.2$ for $m\leq 5$ yields a characteristic correlation length of $\lambda=1.1\pm 0.1$ beads. This short correlation length suggests that entangled chains do not consist of a sequence where every link in the same chain is fully extended, i.e. with the force chain following the links of the chain. Rather, the stretched links can still act as extra constraints on the packing where large forces can be supported, but these forces may be transmitted along frictional contacts with other beads instead of just along the links of a single chain.
@@ III.5 Prediction of the stress-strain curve
Now that we have established that strain stiffening occurs along with stretching of links, but only in packings with system-filling clusters of entanglements, we desire a predictive model of how that leads to strong strain stiffening. Since the stiffness of the beads is much higher than that of the elastic boundary, the stiffness of the overall system would be limited by the weaker stiffness of the boundary without constraints that can provide support across the entire system. Importantly, system-filling clusters of entanglements allow for a possible way to provide additional internal constraints that prevent failure by shear banding and provide an effective replacement for confinement at the boundary. With such constraints, the strength can exceed the confining pressure at the boundary by stretching of the chains. While the strength can originate from the stretching of chains, the constraints of entanglement may allow forces to be transmitted along a contact network that includes the entire system when connected by system-filling clusters of entanglements.
To calculate the stress-strain relation we use a mean-field mode where we assume the average stress $\sigma$ measured on the endplate in response to triaxial load ultimately comes from and is equal to a volume average of the magnitude of the stress on a cross-section of particles. This matching of magnitudes regardless of the direction of forces is reasonable if those contact forces are redirected randomly along a force network. To relate the global stress $\sigma$ to the particle stress $\sigma_{p}$ and particle strain $\gamma_{p}$, we write the derivative of the global stress-strain relation using the chain rule from calculus
@@
The angle brackets represent a mean-field average to be calculated as a volume average over a typical local structure. The effective extensional modulus of the chain $\partial\sigma_{p}/\partial\gamma_{p}\equiv E_{p}$ is calculated as the force pulling on the chain divided by the average cross-sectional area of the chain and the strain of the chain (see Appendix B). $\langle\partial\sigma/\partial\sigma_{p}\rangle$ is the ratio of average stress over the entire volume to the average stress on a stretched chain, corresponding to the fraction of the volume that is being stretched, and is calculated as the packing fraction $\phi$ times the probability of stretched links $p_{s}$. $\langle\partial\gamma_{p}/\partial\gamma\rangle$ then relates the particle strain to the applied strain $\gamma$ averaged over links that are stretched. Equation 1 can now be rewritten as
@@
Figure 9: (color online) Fitted curvature $a_{2}$ of the stress-strain data from Fig. 3, including the 30% run-to-run variation in the uncertainty. Dotted line: model prediction for $a_{2}$ due to stretching of links with system-filling clusters of entanglements. The model prediction is consistent with each curvature for $N\geq 9$ where strong strain stiffening and system-filling clusters of entanglements are observed. The vertical line delineates the boundary between weak and strong strain stiffening based on the magnitude of the curvature.
For comparison to the model, a quadratic function $\sigma=a_{2}\gamma^{2}$ was fit to each measured $\sigma(\gamma)$ in Fig. 3 up to $\gamma=0.2$. Fit values of the curvature $a_{2}$ are shown in Fig. 9. For each $N\geq 9$ where strong strain stiffening is found, the predicted curvature $a_{2}$ based on structural measurements with a 25% uncertainty is consistent with each fit including the 30% run-to-run variation on the measurements. The average coefficient of determination for the quadratic fits for $N\geq 9$ is 0.988. Fitting the curve at $N=10000$ up to a larger strain of 0.45 results in less than a 2% change in $a_{2}$, confirming the scaling is similar over a wider parameter range. Since the curvature $a_{2}$ of the measured $\sigma(\gamma)$ for $N\geq 9$ is consistent with the model prediction, the strength of strain stiffening can be attributed to the stretching of links. The quadratic shape of strain stiffening curves is attributed to an assumed linear increase in the fraction of links being stretched $p_{s}$ with strain.
@@ III.6 Scaling behavior
To confirm whether the measured stress-strain curves $\sigma(\gamma)$ follow the predicted quadratic curvature, we replot each $\sigma(\gamma)$ from Fig. 3 on a log-log scale in Fig. 10. The strong strain-stiffening curves with $N\geq 9$ are seen to be parallel to the predicted quadratic curve with slope 2 for strains $\gamma\stackrel{{\scriptstyle>}}{{{}_{\sim}}}0.04$, showing that the quadratic prediction correctly describes these data where strong strain stiffening occurs. A transition from strain softening to mild strain stiffening and then strong strain stiffening can be seen for the curves as $N$ increase, which gradually approach a slope near 2 for large $N$. Thus, the transition from mild to strong strain stiffening is not defined by a sharp change in the scaling law, but rather corresponds to a sharp jump in the magnitude of the curvature of the stress-strain relation between $N=8$ and 9, as shown quantitatively in Fig. 9.
The stress-strain curves in Fig. 10 are nearly linear for small strain $\gamma\stackrel{{\scriptstyle<}}{{{}_{\sim}}}0.01$. This is in a range of stress that is consistently less than 10 kPa, which does not reach significantly higher strength than packings of individual beads ($N=1$, see inset of Fig. 2). Since the strength of unlinked bead packings is understood to be determined mainly by the stiffness of the confining elastic membrane Lambe and Whitman (1969), much of the linear regime for larger $N$ is also expected to be due to the stiffness of the elastic membrane as well. Since this linear range of stress is less than 1% of the stress range that we measure, and our resolution is very limited here, we cannot distinguish the contribution of the elastic membrane from other mechanisms for the linear stress-strain region in the chain packings for $N>1$.
In Eq. 3, we could have assumed a constant added to $p_{s}$ corresponding to the fraction of links fully stretched at zero strain, which could be as large as 0.09 based on the $N=20$ data in Fig. 7b. This would have added a linear term to the stress-strain relation, corresponding to the nearly linear regime in Fig. 10 for small strain. To test if this nearly linear stress-strain regime is related to $p_{s}$ at zero strain, we added a linear term to the polynomial fit $\sigma=a_{1}\gamma+a_{2}\gamma^{2}$ to the data. Fits of the stress-strain relation yield an average $a_{1}=0.3\pm 0.2$ MPa, which would correspond to a probability of stretched links of $p_{s}=a_{1}/E_{p}\phi\langle\partial\gamma_{p}/\partial\gamma\rangle=0.004\pm 0.003$. This is much smaller than the number of links that are fully extended at $\gamma=0$ as seen in Fig. 7, suggesting that most of these initially extended links are not stretched enough for the link stiffness to contribute to the packing strength. Furthermore, much of the contribution to $a_{1}$ is due to the stiffness of the elastic membrane, so the fraction of stretched links at zero strain should be even less than the value inferred from $a_{1}$, and we should not use the value of $a_{1}$ to directly infer a fraction of stretched links. It is reasonable to understand that the fraction of stretched links at zero strain is negligible, as long as the initial packing is done without a significant loading.
@@ IV Generalizing the strain stiffening model to staples
Equation 2 represents a simple model which attributes the large stress in entangled structures to the deformation of particles, with strain stiffening due to an increasing number of deformed particles with strain. This concept could be extended to a variety of entangled particle systems, not just limited to chains. To test the generality of this model for strain stiffening we consider a very different particle shape: rigid U-shaped staples, which can also entangle. Staples have a few features that are helpful for testing the stress-strain model beyond simply having another shape as a data point. Since the particles are rigid, geometric calculations can be made to determine the free entangled area and predict a probability of deformation $p_{s}$ without need for measurements. Because the particles deform plastically, we can measure the bend angles of staple arms after compression of packings to infer information about entanglement and extensional forces between particles without the need for x-ray tomography, and thus obtain data more easily, which is helpful for example to test the strain dependence of factors such as $p_{s}$ and $\langle\partial\gamma_{p}/\partial\gamma\rangle$.
@@ IV.1 Strain stiffening
Figure 11: (color online) Another example of strain stiffening in randomly packed particles that can entangle – in this case rigid U-shaped staples. The black curve indicates the mean and the gray band indicates the standard deviation of three measurements. The red dotted line shows the model prediction of Eq. 6 due to stretching of staples, which is consistent with the measured curvature within the 37% uncertainty on the model. The blue dashed line shows the best quadratic fit.
@@
The staples enclose an area that corresponds to a free enclosed area $A$ of 80 times the arm or backbone cross-section $wt$ (see Fig. 1c), which is large compared to the longest chains we measured in Fig. 6, so a large number of entanglements is expected. Likewise, simulations of staple packings for the aspect ratio $L_{x}/L_{y}=0.37$ produced an average of 6 entanglements per particle Gravish et al. (2012), more than enough for system-filling clusters according to the Erdös-Rényi model. Our observation of strain-stiffening staple packings is consistent with our observation for chains in Sec. III.2 that average effective free enclosed areas that are large enough to entangle more than one particle result to strain-stiffening behavior. Since the number of entanglements as a function of arm length has already been studied Gravish et al. (2012), here we focus on identifying the source of stress and extending the model for the stress-strain relation to staples, which is less straightforward.
@@ IV.2 Particle bending observations
To identify the source of stress, we look at plastically deformed staples after triaxial compression of a packing. Figure 12 shows a random sampling of 102 plastically bent staples that were poured out and sorted after a triaxial compression test with a maximum strain of 47%. The bends can be categorized based on where contact forces would have to be to produce the bends. 20% of staples showed no resolvable plastic deformation, indicating that the stresses on those particles were below the yield stress for plastic deformation. 45% of staples have the arms bent outward, suggesting they were pulled apart due to extensional loadings on both arms. Another 19% of staples had arms bent outward with an additional inward bend in the backbone. This bend of the backbone could be due to bending from tensile forces on the arms if the contact points are offset from the backbone, or the backbone could be bent due to an additional contact pushing directly on the backbone along with tensional forces on each arm. Combined, this is a total of 64% of the staples that are clearly under some sort of tension. 17% of staples had deformations that indicate a mixture of extensional forces on one arm and compression loadings on the other arm, corresponding to a more complex force balance requiring multiple contact points for mechanical equilibrium which are not apparent from the plastic deformation alone. Notably, no staples were bent in such a way to indicate compressive forces on both arms, even though the sample overall was under compression. Overall, the vast majority of deformed staples experienced extensional forces on the arms, indicating extensional forces from entangling were dominant. This is consistent with previous simulations of packings that found tensile stress on S-shaped and U-shaped particles Karapiperis et al. (2022), as well as expectations based on the observations of chains in Sec. III.4 that tension on particles plays an important role in strain stiffening.
@@ IV.3 Model for stress
The model of Eq. 2 for the strength of entangled packings expresses the slope of the stress-strain curve in terms of the probability of particles being stretched $p_{s}$, and the ratio of local strain to global strain $\langle\partial\gamma_{p}/\partial\gamma\rangle$. Since the plastic deformation of staples is more easily expressed in terms of a permanent plastic deformation angle of the bent arms $\theta_{p}$ instead of the linear strain of a chain $\gamma_{p}$, we rewrite the chain rule of Eq. 1 in terms of the arm bend angle $\theta_{p}$. While the load on particles is applied in the direction of tension, the bending in Fig. 12 suggests that the internal stress is mainly bending stress at the corners of the staples, and perhaps in the backbone. For extensional forces $F_{p}$ applied to the arms at a lever arm distance $l$ away from the corners of the staple, the average stress at the surface of the staples due to bending is approximately $\sigma_{p}=6lF_{p}L_{y}/wt^{2}(L_{y}+2L_{x})$ (see Appendix F for derivation of relationship between normal stress and force). Thus, an appropriate application of Eq. 1 and 2 to staples is